OpenStudy (anonymous):

Describe the end behavior of f(x)=-3x^36+7x^3-12x+13

OpenStudy (anonymous):

degree is even (it is 36) leading coefficient is negative ( it is -3) so goes to \(-\infty\) as \(x\to \infty\) and also as \(x\to -\infty)\)

OpenStudy (anonymous):

thank you:)